15 April 2012 UT Science Prep checkout

Snow and high winds. Ice on the dome and weather mast which precludes opening. Power outages in the morning, after which they went to generator power. Returned to line around 5pm. But the weather is not good. Steve Allanson talked to Rick and they are leaving MODS off, until it will be used for observing. Doing closed-dome LUCI calibrations and basic LBC checkout.

LUCI calibrations:
  • Problem putting in cal unit via the web-IO interface. Neither as observer nor as engineer could we login. Access denied - not in configuration mode But this was cleared (although we don't know why).
  • Got some error messages about WCS keywords. These messages did not appear when running the calibration scripts (below), but we do not know why (GEIRS was restarted, GEIRS Server restarted, and Telescope Service was restarted).
 
Getting data from telescope service
Unable to get WCS keywords!
Error assigning WCS keywords.
  • Problem running scripts. Database is corrupt again. Does every power outage corrupt the instrument database?!
    • Commented out INSTRUMENT section of scripts and these ran successfully.
  • Tables of focus offsets are missing
  • Ran zJ calibrations.

LBC basic checkout:
  • Step 1: send preset in closed dome. Slews telescope but fails to move blue filter.
    • Blue filter wheel motion hung. Blue HK is not reporting any values. Blue HK has dropped out during the past days.
    • Errors in log reported in IT #4050 (LBC Blue connection with PortServer flaky (HK and filter connection problems)).

06:15 UT.
  • Steve parking telescope for the rest of the night.
  • LUCI closed-dome checks done.
    • need on-sky time to:
      • determine transformation matrix (Doug, John)
      • check offsets
    • Problems:
      • database corrupted so we cannot run scripts (must comment out instrument setup to run script).
      • filter focus tables not present.
      • errors with logging into web-io to move cal unit & WCS keywords (mysteriously) resolved.
  • LBC problems with portserver communication (192.168.59.114). IT #4050.

-- OlgaKuhn - 15 Apr 2012
Topic revision: r2 - 15 Apr 2012, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback